What is the 30th term of the sequence: -34, -29, -24, -19
Answer by Edwin McCravy(20056)

Instead of doing your problem for you, I will instead do
one just exactly in every detail step-by-step like yours:

What is the 38th term of the sequence: -35, -31, -27, -23
2nd term - 1st term = (-31)-(-35) = -31+35 = 4
3rd term - 2nd term = (-27)-(-31) = -27+31 = 4
4th term - 3rd term = (-23)-(-27) = -23+27 = 4

Since those are all the same the sequence is arithmetic
with common difference d = 4 and first term a1 = -35,
and number of term n = 38

So substitute a1 = -35, n = 38, and d = 4 in

an = a1 + (n-1)d
a30 = -35 + (38-1)4
a30 = -35 + (37)4
a30 = -35 + 148
a30 = 113

No do yours exactly the same way, step by step.

Edwin